OpenStudy (anonymous):

Which of the following is the most simplified version of ^3{sqrt(432)} ?

OpenStudy (anonymous):

@midhun.madhu1987

OpenStudy (anonymous):

@Lyssa123

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@camerondoherty

OpenStudy (anonymous):

it looks like 

OpenStudy (solomonzelman):

432 = 2× 2³ × 3³ =2 × (2 × 3)³ = 2 × 6³

OpenStudy (solomonzelman):

So how can you simplify it ?

OpenStudy (anonymous):

i dont know

OpenStudy (solomonzelman):

-:(

OpenStudy (midhun.madhu1987):